"An optical pump of wavelength λ1 nm is used for excitation of a laser of wavelength λ2 nm. If the pump has a power of X W, and if one in every 10 pump photons is absorbed by the laser crystal for the emission process, calculate the peak power of the laser and number of laser photons emitted per second
"
λ2 = 543.75 W = 13.84
1
Expert's answer
2020-05-21T13:10:59-0400

As per the question,

wavelength of the photon "(\lambda)=543.75nm"

Pulse width "=13.84nm"

Peak power"=\dfrac{E}{w}=\dfrac{hc}{\lambda \times 13.84 \times 10}"


"=\dfrac{6.6\times 10^{-34}\times 3\times 10^8}{543.75\times 10^{-9}\times 13.84\times 10^{-9}\times 10}"

"=2.64\times 10^4watt"

Number of photon (n)"=\dfrac{P}{E}=2.64\times 10^3=2640"